OpenStudy (anonymous):

Solve for x (4/x) + (4/x^2 - 9) = (3 / x - 3) Solve for x (4/x) + (4/x^2 - 9) = (3 / x - 3) @Mathematics

OpenStudy (anonymous):

x = 4 and x = 9 x = –4 and x = 9 x = 4 and x = –9 x = –4 and x = –9

OpenStudy (anonymous):

Those are the options

OpenStudy (anonymous):

x=(1/12)(1-sqrt(97)) or x=(1/12)(1+sqrt(97))

OpenStudy (anonymous):

None of the options had that answer

OpenStudy (anonymous):

its -4 and 9 I copied the question wrong

OpenStudy (anonymous):

Oh ok thank you :) let me see if I can get that exact answer

OpenStudy (anonymous):

Yeah I got it thank you :)

OpenStudy (anonymous):

any time